# Flexible Multitask Learning with Factorized Diffusion Policy

### v1

@@ IV-C Multitask Learning
Real-world. We further evaluate our method in real-world settings on two tasks: cube red (300 demonstrations) and hang low (200 demonstrations). 20 samples are used for evaluation, and results are summarized in Table III. The DP baseline often overfits to specific joint trajectories, failing to attend to RGB inputs due to the multimodal and perceptually complex nature of the tasks. By contrast, FDP captures diverse behavior patterns more effectively by decomposing the action distribution across sub-modules. This results in higher success rates. Fig. 3 shows qualitative failure cases from baseline methods, which struggle to capture the complex distribution, resulting in imprecise end-effector poses and frequent task failures.
@@ IV-D Task Transfer and Adaptation
In this section, we evaluate the adaptability of FDP in adapting to novel tasks under limited data. We compare several adaptation strategies: full-parameter fine-tuning, partial fine-tuning of the router, observation encoder, and selective module expansion via new expert components. The proportion of tunable parameters of FDP under different adaptation strategies are (a) router-only activates 0.5%, (b)+ observation encoder activates 11%, and (c)+ new module activates 27% of parameters.
Method Policy Cube Blue Hang High Avg.
Full Param. DP 0.750 0.850 0.800
SDP 0.700 0.800 0.750
MoDE 0.750 0.800 0.775
FDP 0.850 0.750 0.800
Router + Obs. Enc. SDP 0.500 0.450 0.475
MoDE 0.500 0.550 0.525
FDP 0.550 0.550 0.550
+ New Module SDP 0.650 0.550 0.600
MoDE 0.700 0.650 0.675
FDP 0.850 0.850 0.850
@@
Simulation. We evaluate adaptation performance on 4 MetaWorld tasks and 4 RLBench tasks, using 10 and 25 demonstrations per task, respectively. We run 60 evaluations for MetaWorld and 40 evaluations for RLBench. As shown in Table II, full-parameter fine-tuning achieves strong performance but is computationally intensive. Partial fine-tuning – modifying only the router or including the observation encoder – offers limited gains. In contrast, adding new modules (two expert blocks per layer for MoE-based methods and a new diffusion component for FDP) consistently improves performance. FDP benefits most from this strategy, leveraging its compositional structure to reuse prior knowledge while efficiently learning new behaviors.
Real-world. We further evaluate adaptation on two real-world tasks, each with 100 demonstrations. 20 samples are used for evaluation. Results in Table IV echo the simulation trends. While full-parameter fine-tuning performs reasonably well, it is resource-intensive. Partial fine-tuning yields modest improvements. The most effective strategy across all methods involves introducing new modules. Under this setting, FDP achieves the best performance, highlighting the advantage of its modular design for rapid and robust adaptation even in complex, real-world scenarios.
@@
TABLE V: Multitask performance of FDP with different numbers of components. Performance improves up to 4 components and plateaus thereafter. We report mean and standard error over 5 seeds
IV-E Analysis
IV-E1 Scaling of Number of Diffusion Components
@@
We view the number of components as a policy-level hyperparameter, analogous to the number of transformer layers or experts in other architectures. Intuitively, if we consider a latent space of observation–action pairs from all demonstrations, the ideal number of components would correspond to the number of clusters in this space, with each component modeling a coherent subset of behaviors [liu2023unsupervisedcompositionalconceptsdiscovery]. In practice, we find that 4–6 components provide a good trade-off between model complexity and performance for the tasks considered. While inference cost scales linearly with the number of components, FDP supports deployment optimizations such as pruning (see Section IV-E4), distillation, or merging [biggs2024diffusionsoupmodelmerging] to reduce overhead.
IV-E2 Scaling of Number of Tasks
@@
IV-E3 Scaling of Number of Demonstrations
@@
Multitask Learning. We evaluate how FDP benefits from increasing amounts of demonstration data. As shown in Fig. 5(a), performance improves steadily with more demonstrations. FDP consistently outperforms baselines, with particularly large gains on RLBench where complex, contact-rich interactions make effective decomposition especially valuable.
Task Adaptation. We analyze how adaptation performance scales with the number of demonstrations, and compare the proposed + New Module strategy with full-parameter fine-tuning. As shown in Fig. 5(b), both strategies benefit from more data, but + New Module achieves comparable or better performance with even fewer demonstrations (on RLBench). This highlights the strength of our modular design in enabling data-efficient adaptation while avoiding the cost and potential catastrophic forgetting associated with updating all model parameters.
IV-E4 Partial Reconstruction of Action Distribution
@@ IV-E6 Knowledge Retention
We investigate the model’s ability to retain knowledge from base tasks during adaptation, a critical property for scalable lifelong learning systems. Conceptually, since the pretrained diffusion experts in FDP remain frozen, the core motor skills are inherently preserved. While a distribution shift in the observation encoder (which is trained from scratch in this work) or a reallocation of weights by the router can occur, these effects can be mitigated by employing a frozen vision foundation model like CLIP [radford2021clip] or by caching the minimal router checkpoints. More generally, to support FDP as a lifelong learning system, we evaluate a strategy using a small replay buffer containing 5 demonstrations per pretraining task. As shown in Table VIII, while adaptation without a buffer leads to some performance degradation, the inclusion of a minimal buffer allows the model to retain nearly all of its original performance. These results suggest that factorization provides a robust foundation for knowledge retention, with further investigation into long-term lifelong learning dynamics reserved for future work.
IV-E7 Diffusion Components Analysis
@@
FDP’s structure contrasts with baseline MoE-based policies. In MoDE, experts specialize according to diffusion noise levels rather than task semantics, leading to noise-level specialization that lacks behavioral interpretability. In SDP, sub-skills emerge from sets of experts selected across layers, making it difficult to assign functionality to any single expert. Experts can be reused across different combinations or ignored altogether. Furthermore, SDP routers tend to favor a small subset of experts, leading to poor load balancing and limited diversity [wang2024sparse_dp_moe]. FDP assigns each behavioral mode to a distinct, standalone diffusion component. This avoids routing instability or expert redundancy commonly seen in MoE. This modular design facilitates straightforward analysis and reuse, contributing to better training stability and more coherent specialization.
IV-E8 Training Convergence of Policies

### v2

@@ IV-C1 Simulation
IV-C2 Real-world
We further evaluate our method in real-world settings on two tasks: cube red (300 demonstrations) and hang low (200 demonstrations). 20 samples are used for evaluation, and results are summarized in Table III. The DP baseline often overfits to specific joint trajectories, failing to attend to RGB inputs due to the multimodal and perceptually complex nature of the tasks. By contrast, FDP captures diverse behavior patterns more effectively by decomposing the action distribution across sub-modules. This results in higher success rates. Fig. 3 shows qualitative failure cases from baseline methods, which struggle to capture the complex distribution, resulting in imprecise end-effector poses and frequent task failures.
@@ IV-D Task Transfer and Adaptation
In this section, we evaluate the adaptability of FDP in adapting to novel tasks under limited data. We first pretrain policies on the tasks shown in Table I, then evaluate adaptation performance on new tasks shown in Table II. We compare several adaptation strategies: full-parameter fine-tuning, partial fine-tuning of the router, observation encoder, and selective module expansion via new expert components. The proportion of tunable parameters of FDP under different adaptation strategies are (a) router-only activates 0.5%, (b)+ observation encoder activates 11%, and (c)+ new module activates 27% of parameters.
Method Policy Cube Blue Hang High
Full Param. DP 15 / 20 17 / 20
SDP 14 / 20 16 / 20
MoDE 15 / 20 16 / 20
FDP 17 / 20 15 / 20
Router + Obs. Enc. SDP 10 / 20 09 / 20
MoDE 10 / 20 11 / 20
FDP 11 / 20 11 / 20
+ New Module SDP 13 / 20 11 / 20
MoDE 14 / 20 13 / 20
FDP 17 / 20 17 / 20
@@
IV-D1 Simulation
We evaluate adaptation performance on 4 MetaWorld tasks and 4 RLBench tasks, using 10 and 25 demonstrations per task, respectively. We run 60 evaluations for MetaWorld and 40 evaluations for RLBench. As shown in Table II, full-parameter fine-tuning achieves strong performance but is computationally intensive. Partial fine-tuning – modifying only the router or including the observation encoder – offers limited gains. In contrast, adding new modules (two expert blocks per layer for MoE-based methods and a new diffusion component for FDP) consistently improves performance. FDP benefits most from this strategy, leveraging its compositional structure to reuse prior knowledge while efficiently learning new behaviors.
IV-D2 Real-world
We further evaluate adaptation on two real-world tasks, each with 100 demonstrations. 20 samples are used for evaluation. Results in Table IV echo the simulation trends. While full-parameter fine-tuning performs reasonably well, it is resource-intensive. Partial fine-tuning yields modest improvements. The most effective strategy across all methods involves introducing new modules. Under this setting, FDP achieves the best performance, highlighting the advantage of its modular design for rapid and robust adaptation even in complex, real-world scenarios.
@@ IV-D2 Real-world
TABLE V: Multitask performance of FDP with different numbers of components. Performance improves up to 4 components and plateaus thereafter. We report mean and standard error over 5 seeds.
V Analysis
V-A Scaling of Number of Diffusion Components
@@
We view the number of components as a policy-level hyperparameter, analogous to the number of transformer layers or experts in other architectures. Intuitively, if we consider a latent space of observation–action pairs from all demonstrations, the ideal number of components would correspond to the number of clusters in this space, with each component modeling a coherent subset of behaviors [24]. In practice, we find that 4–6 components provide a good trade-off between model complexity and performance for the tasks considered. While inference cost scales linearly with the number of components, FDP supports deployment optimizations such as pruning (see Section V-D), distillation, or merging [4] to reduce overhead.
V-B Scaling of Number of Tasks
@@
V-C Scaling of Number of Demonstrations
@@
V-C1 Multitask Learning
We evaluate how FDP benefits from increasing amounts of demonstration data. As shown in Fig. 5(a), performance improves steadily with more demonstrations. FDP consistently outperforms baselines, with particularly large gains on RLBench where complex, contact-rich interactions make effective decomposition especially valuable.
V-C2 Task Adaptation
We analyze how adaptation performance scales with the number of demonstrations, and compare the proposed + New Module strategy with full-parameter fine-tuning. As shown in Fig. 5(b), both strategies benefit from more data, but + New Module achieves comparable or better performance with even fewer demonstrations (on RLBench). This highlights the strength of our modular design in enabling data-efficient adaptation while avoiding the cost and potential catastrophic forgetting associated with updating all model parameters.
V-D Partial Reconstruction of Action Distribution
@@ V-F Knowledge Retention
We investigate the model’s ability to retain knowledge from base tasks during adaptation, a critical property for scalable lifelong learning systems. Conceptually, since the pretrained diffusion experts in FDP remain frozen, the core motor skills are inherently preserved. While a distribution shift in the observation encoder (which is trained from scratch in this work) or a reallocation of weights by the router can occur, these effects can be mitigated by employing a frozen vision foundation model like CLIP [27] or by caching the minimal router checkpoints. More generally, to support FDP as a lifelong learning system, we evaluate a strategy using a small replay buffer containing 5 demonstrations per pretraining task. As shown in Table VIII, while adaptation without a buffer leads to some performance degradation, the inclusion of a minimal buffer allows the model to retain nearly all of its original performance. These results suggest that factorization provides a robust foundation for knowledge retention, with further investigation into long-term lifelong learning dynamics reserved for future work.
V-G Diffusion Components Analysis
@@
FDP’s structure contrasts with baseline MoE-based policies. In MoDE, experts specialize according to diffusion noise levels rather than task semantics, leading to noise-level specialization that lacks behavioral interpretability. In SDP, sub-skills emerge from sets of experts selected across layers, making it difficult to assign functionality to any single expert. Experts can be reused across different combinations or ignored altogether. Furthermore, SDP routers tend to favor a small subset of experts, leading to poor load balancing and limited diversity [39]. FDP assigns each behavioral mode to a distinct, standalone diffusion component. This avoids routing instability or expert redundancy commonly seen in MoE. This modular design facilitates straightforward analysis and reuse, contributing to better training stability and more coherent specialization.
V-H Training Convergence of Policies